OpenStudy (anonymous):

The sum of two positive numbers is 8. Find the numbers if the sum of the cube of one and the square of the other is a minimum.

OpenStudy (anonymous):

this question is so frustratinggggg

OpenStudy (anonymous):

Positive integers?

OpenStudy (anonymous):

yes the answer is meant to be 2 and 6

OpenStudy (anonymous):

Yes that's right.

OpenStudy (anonymous):

y = 8-x, z = x^2+y^3 z = -x^3+25 x^2-192 x+512 min{-x^3+25 x^2-192 x+512} = 44 at x = 6 Y= 2

OpenStudy (anonymous):

you only have a few possibilities, 1 and 7, 2 and 7, 3 and 4, 4 and 4.

OpenStudy (anonymous):

And you need a low number for the cube.

OpenStudy (anonymous):

So you need only look at the first two.

OpenStudy (anonymous):

2 and 6 I mean.

OpenStudy (anonymous):

Typos:-( you only have a few possibilities, 1 and 7, 2 and 6, 3 and 5, 4 and 4

OpenStudy (anonymous):

Thanks guys !